### v2

# Words as Trigger Points in Social Media Discussions—A Large-Scale Case Study about UK Politics on Reddit

Abstract.
Trigger points are a concept introduced by Mau et al. (2023) to study qualitative focus group interviews and understand polarisation in Germany. When people communicate, trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned. In the original studies, individuals show strong and negative emotional responses when certain triggering words or topics are mentioned. In this paper, we introduce the first systematic study of the large-scale effect of individual words as trigger points by analyzing a large amount of social media posts. We examine online deliberations on Reddit between 2020 and 2022 and collect >>>100 million posts from subreddits related to a set of words identified as trigger points in UK politics. We find that such trigger words affect user engagement and have noticeable consequences how online discussions occur. Trigger words cause animosity and incentivise hate speech, adversarial debates, and disagreements. Our work is the first to introduce trigger points to computational studies of online communication. The findings are relevant to researchers interested in affective computing, online deliberation, and those examining how citizens debate politics and society in light of affective polarisation.
Affective Computing, Polarization, Reddit, Trigger Words, User Engagement
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: THE WEB CONFERENCE 2025; 28 April–2 May, 2025; Sydney, Australiaccs: Information systems World Wide Webccs: Computing methodologies Natural language processing
1. Introduction
Trigger points, introduced by Mau et al. (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions. Examining trigger points in online discussions helps understand why and when social media users engage in disagreements or affective political deliberations. This opens the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
Figure 1. Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
Trigger points are related to how individuals perceive themselves in society. People have subjective expectations about acceptable behaviour, such as explicit knowledge of law-like rules and, more importantly, an implicit understanding of appropriate actions. These deeply running beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world. People feel triggered if these normative understandings are questioned. More formally, Mau et al. (2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract. Trigger points jeopardise what individuals consider as the fabric of society and their role in it.” Trigger points can affect individuals through four mechanisms.
  • Inequality. When expectations about equality might not be met. Either similar people are treated unequally, or, to the contrary, people who should be treated differently are treated the same. For example, social benefits might be at eye level with minimal wage earners.
  • Norm transgression. It is the transgression of what is considered “normal” behaviour. A typical example relates to the lavish lifestyle of the privileged.
  • Slippery slope. There could be a fear of society developing norms about appropriate behaviour that goes in subjectively undesired directions and the impression that the individual might not have the power to do something against it. “Opening the borders so that a country could be flooded with immigrants” is an example in this context.
  • Behavioural demands. When society may be perceived as demanding unreasonable behaviour. The role of pronouns is a point in case where the generic masculine pronouns “he/him” might trigger some, and the neutral pronoun “they/them” might trigger others.
Going against deeply held beliefs about individuals and their societal role causes discomfort. People respond to such triggers in an affective behavioural mode and increase a debate’s intensity and emotionality.
In the social media context, we expect that this has two direct consequences (Figure 1). Trigger points lead to:
  • Hypothesis 1 more user engagement, which incites a higher number of messages or comments;
  • Hypothesis 2 higher levels of animosity, which incite polarisation, negativity, anger, and hate.
Our study focuses on how individual words can act as trigger points in online debates. We look into five words (Rwanda, Brexit, NHS, Vaccine, Feminism) we expect to work as trigger points in the context of UK politics. Collecting >>>100 million related Reddit posts, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses. The analysis builds upon a systematic comparison between a treatment and control group in three different scenarios:
  1. (1)
    Space: specific subreddits where the selected words trigger users.
  2. (2)
    Semantics: words similar in meaning to the target words but without the triggering component.
  3. (3)
    Time: time periods identified as relevant for each specific trigger word.
Our results show that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
2. Related Work
Trigger points (Mau et al., 2023) bear great potential to understand polarisation and division in online discussions. However, the existence or validity of the concept has not been tested on a large scale. Trigger points can be particularly relevant to work on social media. They help understand the causes of disagreements in online conversations (De Kock and Vlachos, 2021) and reasons for their escalation (De Kock and Vlachos, 2022; van der Meer et al., 2023). Further to triggering ad hominem attacks, escalation can lead to hate speech. There is, however, no unanimous agreement on how it differs from abusive and offensive language (Basile et al., 2019; Davidson et al., 2017; Talat and Hovy, 2016).
Despite the rising interest in hate speech and harmful language in social media platforms, including its implicit forms (Ousidhoum et al., 2019; Sap et al., 2019; ElSherief et al., 2021; Ocampo et al., 2023), and dog-whistles which are coded expressions that carry two meanings with a hateful or provocative coded one (Breitholtz and Cooper, 2021), we know little about the sources of these phenomena. Examining what can trigger harmful reactions may help us make more informed choices when collecting data, thus, leading to more robust detection models and mitigating the well-known problem of selection bias (Wiegand et al., 2019; Ousidhoum et al., 2020).
Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al., 2019), Blodgett et al. (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model. Studying trigger points helps track the origins of bias and assess whether a social media post can lead to hateful comments (Dahiya et al., 2021).
Further, understanding how trigger words affect online communication also sheds light on how language affects polarisation on social media (Karjus and Cuskley, 2024). It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon et al., 2022).
3. Methodology
We use the UK context as a large-scale case study to observe polarisation when individuals communicate on social media. Our goal is to report behavioural mechanisms similar to those observed by Mau et al. (2023) when conducting interviews to analyse the perceived polarisation of the German society. Based on a curated set of likely candidates, we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
3.1. Term selection
We selected terms likely to provoke reactions from individuals in the United Kingdom. Specifically, we include National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism in our study. These trigger words were chosen because they are associated with high-profile public debates, policy changes, and social movements, which have been known to provoke significant public engagement and emotional responses. Table 1 indicates the respective mechanism through which the terms may trigger individuals.
Rwanda Brexit NHS Vaccine Feminism
Inequality
Norm transgression
Slippery Slope
Behavioural Demands
Table 1. Trigger words and their respective trigger mechanisms (described in the introduction).
Rwanda.
Rwanda plays a particular role in the UK’s immigration policy. [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible Until late 2021, Rwanda was just the name of any other African country. However, in early 2022 the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership. Instead of processing asylum cases directly in the UK, migrants would be flown to Kigali, and their claims were to be assessed there, in turn for U\$300m, (2.3% of Rwandan GDP) and expenses for every migrant sent. With this, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist, 2023).
Brexit.
The UK’s departure from the European Union has been controversial since its inception. Even eight years post-exit, debates persist regarding the soundness of the decision. The division among the population was already evident during the tightly contested 2016 referendum, which revealed distinct trends among various demographic groups based on location, age, and income (Arnorsson and Zoega, 2018). Since the referendum, further divisions have emerged concerning the future relationship between the UK and EU (Vasilopoulou and Talving, 2019). Additionally, misinformation and disinformation have intensified these debates (Höller, 2021; Bruno et al., 2022). Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from the East of Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
NHS.
Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the health care system (The Guardian, 2022; The Independent, 2024). As a result, the NHS has become a hot topic of discussion among the UK population. Public opinion is divided, with some supporting that it receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick et al., 2015). This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
Vaccine.
Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years. Even after their widespread roll-out, discussions over their effectiveness and safety are ongoing. Initial reluctance was observed across various demographic segments, influenced by factors such as education level, political affiliation, and socioeconomic status (Troiano and Nardi, 2021). Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell, 2021). Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al., 2022; Garett and Young, 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
Threads Comments Year
Rwanda 0.05 1.87 2020, 2022
Brexit 0.37 11.57 2020 - 2022
NHS 0.34 6.20 2020 - 2022
Vaccine 1.42 28.11 2019, 2022
Feminism 0.94 17.79 2020 - 2022
EU 0.2 12.4 2020 - 2022
Red Cross 0.02 0.2 2020 - 2022
SubSaharan 0.08 1.5 2020, 2022
Surgery 0.5 3.9 2019, 2022
Woman 2.1 19.3 2020 - 2022
Table 2. Numbers, in millions, of threads, comments, and time period for each trigger word.
Feminism.
Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz, 2019). Previous work has shown that women politicians and journalists tend to be harassed online every few seconds (Posetti et al., 2022).  [^2] https://www.amnesty.org.uk/press-releases/women- abused-twitter-every-30-seconds-new-study Some social media influencers have built a large audience by sharing and promoting harmful content against women. This has led to the formation of online echo chambers (Farrell et al., 2019) where such views are amplified and propagated. Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope and unreasonable behavioural demands.
Keyword Subreddits
Feminism
AskMen, MensRights,
antifeminists, AndrewTateTalk
Brexit
neoliberal, brexit, ireland, unitedkingdom,
The_Donald, ukpolitics, tories,
NHS
neoliberal, CoronavirusUK, unitedkingdom,
ukpolitics, tories, labourUK
Rwanda
neoliberal, tories, labourUK, currenteventsuk
ukpolitics, uknews, unitedkingdom
Vaccine
conspiracy, neoliberal, unpopularopinion,
antifascistsofreddit, coronavirusuncensored,
CoronavirusUK, Coronavirus, wuhan_flu
Table 3. Subreddits selected for each trigger word.
Feminist Rwanda Brexit NHS Vaccine
Be Af DiD CI Be Af DiD CI Be Af DiD CI Be Af DiD CI Be Af DiD CI
Treatment
Only
47 53 45 55 46 54 46 54 47 53
Space 49 51 4.1 (3.8, 4.5) 50 50 9.7 (8.2, 11.3) 49 51 6.3 (6.1, 6.6) 49 51 6.1 (5.7, 6.4) 49 51 4.1 (3.5, 3.9)
Semantic 48 52 4.6 (3.9, 5.2) 50 50 9.8 (6.6, 14.9) 47 53 3.2 (2.6, 3.7) 53 47 13.4 (8.7, 18.1) 49 51 5.0 (3.1, 6.9)
Temporal - 49 51 7.2 (2.9, 11.6) - - 50 50 5.9 (4.1, 7.6)
Table 4. Percentage of comments per thread before and after each trigger appearance (Hypothesis 1). Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Semantics, comparisons. Bold scores indicate statistically significant results. Detailed distribution of comments can be seen in Appendix B, see Figure 7.
3.2. Data collection
Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al., 2020). We searched for exact matches of specific keywords in comments from 2019 to 2022, and then collected the entire discussion thread for each extracted comment [^3] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume.. We excluded posts in languages other than English, as identified with fastText-based language identifier (Bojanowski et al., 2017). Posts marked as “[deleted]” by the API were also removed. Following this, only threads that contained posts before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
As a final filtering step we limited our analysis to posts occurring 30 minutes before and 30 minutes after the first mention of the triggering word. Previous research has shown that Reddit posts typically receive the majority of interactions within a short window after the post is created (Thukral et al., 2018). Moreover, early responses are often important in shaping the direction of the conversation (Singer et al., 2016; Horne et al., 2017). Considering the fast-paced nature of Reddit, the 30 minutes window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data. Table 2 shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered.
3.3. Research design
We identify the causal effect of trigger points with the difference-in-difference estimator (DiD) which systematically compares a treatment with a control group (Angrist and Pischke, 2009; Card and Krueger, 1994). It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group. In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words. There should be a more substantive behavioural change for the treatment group than for the control group.
3.3.1. Control groups
We systematically compare treatment and control groups along three different dimensions. [^4] Identifying the causal effect in three completely unrelated comparisons helps address concerns about eventual endogeneity.
1. Space
Discussions on Reddit are structured in different fora, also called subreddits. Aiming to investigate the different effects a trigger word might have on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 3). We compare them with discussions in a control group, i.e., the rest of Reddit with many non-domain specific subreddits.
2. Semantics
We also compare trigger words with a term that shares a similar meaning, but that does not go against the deeper underlying transgressions responsible for trigger responses. To members of the subreddits in Table 3, using an explicit trigger word or its more harmless control term should make a real difference. We define related control group words and contrast woman to Feminism/Feminist, Red Cross to NHS, EU/European Union to Brexit, surgery to vaccine, and names of Rwanda’s neighbouring countries [^5] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia. to Rwanda.
3. Time
In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time. For “Rwanda” we compare discussions in subreddits likely to incite reactions in 2020 to discussions in 2022. For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic). Again, we expect that users from the subreddits in Table 3 are triggered by those only in 2022 but not in the respective earlier years.
3.3.2. Statistical analysis
We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 4) and 48 experiments for H2 (Table 5). For each experiment, we prepare the data in the same way. The treatment/placebo is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread. To test H1, we count the number of posts for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group. We then calculate the share of posts before and after the intervention. For example, if there are 8 posts during the 30 minutes before the trigger word and 12 posts during the 30 minutes after the trigger word, we normalise the data of this discussion thread to 40% before intervention and 60% after intervention. For experiments with H2, we count the number of the respective animosity in a thread and then normalise the data as a share of animosity per thread before and after the trigger word.
We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke, 2009):
[…]
where proportion is the per-thread-normalised target feature and the interaction term is given by
$\text{interaction}=\text{trigger\_dummy}\times\text{treatment\_dummy}.interaction = trigger_dummy × treatment_dummy$ .
Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al., 2023). It requires that in the absence of treatment, the difference between the outcomes of the treated and control groups would follow the same trajectory over time. We investigate the plausibility of the assumption in section 4.3.
3.4. Features
We rely on several features to examine our initial hypotheses and study the impact of trigger words in online discussions. We simply use post counts for the first hypothesis (H1) related to user engagement. The second hypothesis (H2) requires a measure of animosity. To identify relevant features that can operationalise animosity, we build on various language models that follow the RoBERTa architecture (Liu et al., 2019) and are trained on social media corpora, specifically X (Twitter). Even though posts on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text. Hence, we expect the trained models to perform well in either context (Priya et al., 2019).
3.4.1. User engagement (Hypothesis 1)
Our first Hypothesis suggests that trigger words lead to higher user engagement. We simply rely on the number of posts to investigate this. That is, we expect to see an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
3.4.2. Animosity (Hypothesis 2)
Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalise this with controversial comments, the presence of negative sentiment, anger, and different forms of hate speech.
Controversiality
We examine whether a comment is controversial using the Reddit controversiality feature. Reddit offers a measure that identifies a comment as controversial if it receives similar amounts of upvotes and downvotes. This metric is based on the users’ approval, so care must be taken when considering already controversial subreddits. We expect trigger words to cause more controversiality.
Figure 2. Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
Sentiment analysis
For sentiment analysis, we rely on the model twitter-roberta-base-sentiment-latest (Loureiro et al., 2022) which identifies a post as negative, neutral, or positive. This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal et al., 2019). The sentiment in social media content has often been used to gain insights into the information related to controversial topics such as vaccination (Melton et al., 2021), politics (Guimaraes et al., 2019), and gender biases (Marjanovic et al., 2022). We expect trigger words to cause more negative sentiments.
Negative emotions
We rely on twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al., 2022) to assign one or more emotions to each tweet. This model is trained with data from the Semeval 2018 Affect in Tweets task (Mohammad et al., 2018), covering 11 different emotions [^6] Detailed list of emotions can be found in Appendix A.1). For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
Figure 3. Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
Hate speech
Finally, we use the twitter-roberta-base-hate-multiclass model (Antypas and Camacho-Collados, 2023) to detect hate speech. It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups [^7] Detailed list of categories can be found in Appendix A.2). The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al., 2021), women’s rights (Farrell et al., 2019), and health issues (Walter et al., 2022). Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist, racism for Rwanda, while we consider all hateful comments jointly for NHS, Brexit and Vaccine.
4. Results
In this section, we present the results for the two hypotheses tested in this paper.
4.1. User engagement (Hypothesis 1)
Our initial hypothesis is that if trigger words significantly influence online discussions, there should be more messages after mentioning these terms. The DiD estimate reveals significant differences between the frequency of comments in the treatment and control groups for all our trigger words (Table 4).
In the treatment-only subset, the difference between user engagement before and after the trigger word is apparent for all the terms we selected. It is largest with Rwanda where we observe a 10%-points (Table 4) increase. To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group. In every setting, we observe a significant difference in the number of comments in the threads between the treatment and control subsets, with the frequency of comments in the treatment group increasing by a larger factor according to the DiD estimate. The lowest value is a 3.2%-point increase for the semantic comparison in Brexit. The largest increase is for NHS in the semantic control group, where we observe a 13.4% point increase.
When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent. For instance, Figure 3 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment. It shows the overall number of messages across all threads. This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial. Similar patterns are also observed in the semantic and temporal control groups for Rwanda (Appendix B, Figure 9).
Overall, our analysis of the influence of trigger words across different control groups indicates a clear increase in engagement for each trigger word analysed, confirming our initial hypothesis on the effect of trigger words on online discussions.
Trigger Setting Feature DiD CI
Brexit Space Contr 1.0 (0.8, 1.3)
N.Sent 3.3 (3.0, 3.5)
Anger 2.7 (2.4, 3.0)
Hate -0.1 (-0.5, 0.4)
Semantics Contr 1.4 (0.7, 2.1)
N.Sent 2.6 (2.0, 3.1)
Anger 3.8 (2.1, 5.5)
Hate -0.5 (-1.9, 0.9)
Feminist Space Contr 0.3 (-0.2, 0.8)
N.Sent 2.6 (2.2, 3.0)
Anger 1.4 (1.0, 1.9)
Sexism 0.6 (0.1, 1.1)
Semantics Contr 0.1 (-0.8, 1.1)
N.Sent 3.0 (2.3, 3.6)
Anger 1.4 (1.0, 1.9)
Sexism 1.4 (0.5, 2.3)
NHS Space Contr 1.1 (0.7, 1.4)
N.Sent 3.1 (2.8, 3.5)
Anger 2.7 (2.3, 3.1)
Hate -0.3 (-0.9, 0.3)
Semantics Contr 4.5 (-3.2, 12.2)
N.Sent 3.5 (-1.2, 8.3)
Anger 4.2 (-11.6, 20.0)
Hate -1.4 (-12.5, 9.6)
Trigger Setting Feature DiD CI
Rwanda Space Contr 2.3 (0.8, 3.7)
N.Sent 5.5 (4.0, 6.9)
Anger 4.0 (2.7, 5.3)
Racism 1.8 (0.0, 3.6)
Semantics Contr 6.6 (0.2, 13.0)
N.Sent 5.5 (1.5, 9.6)
Anger 5 (-4.6, 14.6)
Racism 14.6 (-0.7, 29.9)
Temporal Contr 5.9 (0.5, 11.3)
N.Sent 4.8 (0.8, 8.8)
Anger 2.8 (-1.0, 6.6)
Racism 0.9 (-11.1, 12.8)
Vaccine Space Contr 1.5 (1.3, 1.8)
N.Sent 3.0 (2.8, 3.2)
Anger 2.0 (1.7, 2.2)
Hate 0.4 (0.0, 0.8)
Semantics Contr 2.2 (1.1, 3.4)
N.Sent 5.4 (4.7, 6.1)
Anger 4.4 (1.6, 7.2)
Hate -0.7 (-2.4, 0.9)
Temporal Contr -0.3 (-3.0, 2.3)
N.Sent 2.9 (1.4, 4.4)
Anger 5.5 (1.4, 9.6)
Hate -2.0 (-5.3, 1.3)
Table 5. Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
4.2. Animosity (Hypothesis 2)
Our second hypothesis suggests that trigger words may increase the frequency of polarising messages and induce animosity in the discussion.
Figure 4. Average controversiality per thread for the treatment and control (space) groups of Rwanda.
Controversiality.
We analyse the average controversiality of each thread and compare the proportion of controversial threads before and after the introduction of the trigger words. Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word. Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the semantic comparison for Rwanda—a causal effect of 6.6% points.
However, significant differences in controversiality do not appear across all settings. Notably, there seems to be no increase in controversial comments when considering Feminism, and in the temporal and semantic settings for Vaccine and NHS, respectively.
Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda for the space experiment, before and after its occurrence. The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points. A similar trend is seen in both the semantic and temporal control settings, as documented in Appendix B, Figure 8).
(a) Rwanda
(b) Feminism
Figure 5. Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
Sentiment & emotion.
We further investigate the effect trigger words have on sentiment and emotion. Based on our hypothesis, trigger words cause users to express more negatively charged emotions. In our analysis, we focus on the presence of negative sentiment and, more narrowly, the presence of anger in emotions.
When considering the negatively charged comments of the treatment and control groups for each trigger word, Table 5 indicates a clear trend. We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the semantics comparison for NHS. The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and semantics experiments.
Let us consider the vaccine trigger, Figure 2, as an example. Here, we again plot the overall number of messages we observe. Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned. However, the treatment group shows more messages than the control group, suggesting that the trigger words affects the expression of negative sentiment.
Similarly to the increase in negative sentiment, trigger words systematically cause more anger. We observe a positive causal effect in all conditions, with only 3 non-significant effects. Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine semantic comparison). As another visual example, the overall number of posts expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure LABEL:fig:emotion_space_vaccine. This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
Hate speech.
As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of hate speech or abusive language for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism. As seen from Table 5, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment. We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and semantic comparison) and Rwanda (space group).
We illustrate our findings again in a visualisation, where we consider all messages irrespective of the thread they are posted in. We observe a distinct pattern, particularly in the context of Rwanda and Feminism, where the volume of hateful content increases after the trigger. Figure 5 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control group in the space experiment.
4.3. Testing the Parallel Trends Assumption.
The DiD estimator relies on the assumption of parallel trends: without the treatment, the treated and control groups would follow the same trajectory over time. While there are no formal guarantees for the assumption to hold, there are means to argue in its favour. The first argument is a theoretical one. In neither of our experimental designs, we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word. Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (semantics), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
We implement event studies for each experiment to further empirically support our argument. The central idea is to model the respective outcome variable as it unfolds over time, which allows us to assess the dynamics of treatment impact before and after the intervention. Plotting the treatment effects at different time points before the treatment, the event study helps visualise whether the pre-treatment trends in the treatment and control groups are indeed parallel. If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al., 2023).
Figure 6. Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
For each of our overall 60 experiments, we divide the 30 minutes before and after the first mention of a trigger word into six distinct time intervals $\{[-30,-20[,[-20,-10[,[-10,0[,[0,10[,[10,20[,[20,30[\}${ [ - 30 , - 20 [ , [ - 20 , - 10 [ , [ - 10 , 0 [ , [ 0 , 10 [ , [ 10 , 20 [ , [ 20 , 30 [ } instead of just two. This allows us to follow, in particular, the pre-treatment trends at a higher granularity. We calculate each discussion thread’s share of (hostile) messages in the six intervals. Averaging over all these data points per period, we subtract the control group from the treatment group and calculate the difference of each observation from the first observation period.
Figure 6 charts the time-wise means over all 60 event studies. If the parallel trends assumption holds, the lines before the treatment at time = 0 should all remain on the horizontal line where the treatment effect = 0. Indeed, this is largely what we find. The mean treatment effect is close to 0 before the trigger word is mentioned for the first time. After the intervention, it is clearly above the horizontal 0-line. [^8] See Appendix for figures that include individual event studies for all experiments from H1 and H2.
5. Conclusion
In this paper, we present the concept of trigger words and study how they affect social media conversations in UK politics. We report empirical evidence based on a large-scale study on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS and Vaccine). In short, our results show that trigger words cause more user engagement, but that this activity is marked by higher levels of animosity.
Our findings confirm the first Hypothesis (H1), which suggests more user interaction after the appearance of a trigger word. We consistently observe more posts across all trigger words and control settings. Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examine four different features that can indicate such effects. Trigger words cause more controversiality, more negative sentiment, and more anger. We also investigate the causal effect of trigger words on hate speech. Where we can, we rely on a model that is capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—. Our findings similarly confirm that trigger words play a causal role in this aspect as well.
6. Limitations
We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset (>>>100 million entries). Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments. Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al., 2023).
Selecting words as trigger points and their corresponding subreddits offers room for improvement. The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community. Similarly, we only investigate a very select set of potential trigger words in the context of the UK.
However, our study had different goal: We consider our work to be a pilot study that intends to investigate whether the trigger points exist in online settings at all. Based on the empirical evidence we find, we can be confident to have found relevant empirical evidence that warrants further investigation.
7. Ethics Statement
All data utilised in our research is sourced from publicly available information or collected using the official Reddit API. All information is provided in an aggregated fashion without reporting sensitive information from individual users.
Appendix A Model Categories
A.1. Emotion Categories
The twitter-roberta-base-emotion-multilabel model classifies each entry in one or more of the following classes: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust.
A.2. Hate Speech Categories.
The twitter-roberta-base-hate-multiclass model classifies each entry in one of the following classes: not_hate, sexism, racism, religion, other, sexual_orientation, disability
Appendix B Extended Results
Figure 7 displays the number of comments posted before and after the introduction of each trigger point on the data collected.
Figure 8 displays the average controversiality per thread before and after the introduction of the Rwanda trigger point in the Temporal and Semantic control settings.
Figure 9 displays the distribution of comments posted before and after the introduction of the Rwanda trigger point in the Temporal and Semantic control settings.
Figure 11 displays the distribution of negative comments, based on their sentiment, posted before and after the introduction of the Brexit, Feminism, Rwanda, and NHS trigger points in the Space control setting.
(a) Brexit
(b) Feminism
(c) NHS
(d) Rwanda
(e) Vaccine
Figure 7. Absolute count of posts before and after each trigger point.
(a) Temporal Setting
(b) Semantic Setting
Figure 8. Average controversiality per thread in the Temporal and Semantic settings and their respective treatment groups, for the trigger point Rwanda.
(a) Semantic Comparison
(b) Temporal Comparison
Figure 9. Rwanda Treatment and Control distributions of comments in the Semantic and Temporal settings.
(a) Brexit
(b) Feminism
(c) Rwanda
(d) NHS
Figure 10. Distribution of negatively charged comments in the Treatment and Control (Space) groups.
Appendix C Extended Test of Parallel Trends Assumption
(a) Experiments from H1.
(b) Experiments from H2.
Figure 11. Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.
Appendix D Computational Resources
An NVIDIA GeForce RTX 4090 GPU was utilised for the experiments conducted. 170 hours spent for the inference process of sentiment analysis, emotion detection, and hate speech detection models.